# Autalic: A Dataset for Anti-AUTistic Ableist Language In Context

### v3

@@ 4.2.3 In-Context Learning
After providing the in-context learning examples, Llama3 (+22.96%) and Gemma2 (+12.68%) display the biggest relative improvement in F-1 scores, indicating that both models benefit from the examples. In particular, we find our ICL examples resulted in large relative improvements in consistency of scores regardless of the language used in the prompts for each LLM. For example, providing Llama with examples helped decrease its relative change in F1 scores from 67.49% to 17.4% when switching from PFL to conceptual prompts, indicating a better understanding of the connection between anti-autistic ableism and ableism toward autistic people.
@@
Figure 4: The mean Cohen’s Kappa scores of each LLM comparing the agreement with human annotators and other LLMs.
4.2.4 Understanding Ableist and Anti-Autistic Speech
@@
Our results with prompt engineering reveal that anti-autistic ableism is too abstract of a concept for LLMs to recognize, providing empirical evidence that their current reasoning abilities are not inclusive of the perspectives of autistic people. LLMs struggle with identifying anti-autistic speech regardless of the terminology used, further indicating they are unreliable agents for data annotation tasks.
@@
Table 3 displays the results of prompt engineering using either person-first or identity-first language and using “anti-autistic” to describe this form of ableism in more conceptual terms. Notably, switching from PFL or IFL to conceptual language in the prompts resulted in the largest relative changes in scores. For instance, prompting Llama with "is this sentence anti-autistic?" instead of "is this sentence ableist toward people with autism?" resulted in a relative increase of 67.49%. These results show that LLMs struggle with understanding that “anti-autistic ableism" and “ableism toward autistic people" refer to the same phenomenon. Even after providing the ICL examples, the relative change in F1 scores between different prompts was as high as 32.66% for Gemma2.
@@
Interestingly, DeepSeek, the only reasoning LLM we test in our study, has the best results and highest consistency out of all the other LLMs. Although its agreement with human annotators is low ($k:0.11$), it is still double that of all other LLMs, as shown in Figure  4. This highlights the difficulty of this task, as more advanced reasoning is required to understand the nuances of anti-autistic ableism, including the terminology we use to describe such speech.
@@ 4.3 Discussion
Classifying anti-autistic ableist speech is challenging even within a Western, English-speaking context, since perceptions of what constitutes anti-autistic content differ significantly even among autistic individuals Keating et al. (2023). Factors such as personal experiences, cultural norms, and evolving discourse around autism advocacy can influence each individual’s perception of or sensitivity toward recognizing toxic speech, making it difficult to establish a consistent classification scheme Ousidhoum et al. (2019); Bottema-Beutel et al. (2021b); Taboas et al. (2023); Kapp et al. (2013). While developing Autalic, we standardize our definition of anti-autistic ableist speech by providing our annotators with an orientation and a glossary. These resources are developed in alignment with the perspectives of autistic people  Bottema-Beutel et al. (2021a); Taboas et al. (2023); Kapp et al. (2013).
Our experiments demonstrate the importance of Autalic in aligning LLM performance to human expectations in the contexts of autism inclusion and ableist speech classification. Through our experiments, we provide empirical evidence of the current limitations of using LLMs and traditional classifiers to identify expressions of anti-autistic ableism. These limitations include: a misalignment with human perspectives, a lack of understanding of the concept of anti-autistic ableism, and a lack of agreement with each other even with in-context learning examples. Each of these limitations adds to the challenge of utilizing LLMs as reliable agents for such tasks.
Standard pre-trained models such as showed poor performance, reinforcing the need for specialized fine-tuning. However, even after fine-tuning BERT on the Autalic dataset, our experiments reveal a high rate of false positives, which can lead to unfair censorship if BERT is employed for this task. Additionally, our results reveal that even state-of-the-art LLMs exhibit low agreement with human annotators on this task, further emphasizing the challenges of detecting subtle forms of ableism using generic models. DeepSeek has the best performance out of all the LLMs in our study, further demonstrating the difficulty of this task, as it is the only a reasoning-focused LLM in our study.
@@ 5 Conclusion
In this paper, we introduced Autalic, the first benchmark dataset focused specifically on the detection of anti-autistic ableist language in context. Through the collection and annotation of 2,400 sentences from Reddit, we aim to fill a critical gap in current NLP research and improve its alignment with neurodiversity.
Looking forward, Autalic paves the way for significant advancements in content moderation systems, hate speech detection models, and research on ableism and neurodiversity. We envision this dataset as a cornerstone for future work in addressing bias against autistic individuals and fostering a more inclusive digital environment. By sharing this resource with the broader research community, we aim to catalyze the development of more equitable NLP systems that better serve underrepresented and marginalized groups.
@@ 6 Limitations
Despite doing our best to include a variety of ableist language against autistic people, our dataset can still show some selection bias Ousidhoum et al. (2020) as we relied on keywords and specific social media threads to collect our data.
We recognize that, like other datasets presented at ACL, our work primarily reflects Western perspectives  Anderson-Chavarria (2022); Kirk et al. (2023); Thapa et al. (2022); Aoyama et al. (2023); Takeshita et al. (2024); Park and Park (2023). Therefore, we do not claim that Autalic is generalizable across languages and cultures as anti-autistic and ableist speech may manifest and be perceived differently. However, we emphasize that Autalic is the first dataset to center the perspectives of autistic people in ableist speech detection, and move away from the tendency of AI to misclassify any text related to disability as ‘toxic’  Narayanan Venkit et al. (2023); Heung et al. (2024); Van Dorpe et al. (2023); Hutchinson et al. (2020). Therefore, it can serve as a reference for future work exploring ableist speech in diverse contexts such as other cultures and languages.
Our dataset contains sentences from Reddit, and subreddits focusing on autism such as r/Aspergers, r/Autism, and r/AutismInWomen were among the most popular subreddits the posts were published in. Our search criteria included words like “r*tard” which is a slur broadly used against many intellectual disabilities and not specifically focused on autism. While we did our best to filter out irrelevant posts, since we broadened our search criteria to include variations of “neurodiver*” such as “neurodiversity”, “neurodiverse”, and “neurodivergent”, the search inevitably yielded posts discussing unrelated topics such as the NEURODIVER video game. The majority of these sentences were also published in 2023. While our dataset is small, the quality of our labels is high due to our standardized annotation protocols and resources that were developed using rigorous and iterative testing.
@@ 7 Ethics Statement
The data collected for our small-scale and non-commerical research is in compliance with Reddit’s API limits and policies  Reddit (2023a, b). All the sentences in our dataset are publicly available, and we follow the methodologies of prior work in our data collection process  Atuhurra and Kamigaito (2024).
We specifically recruit annotators for whom this collaboration and resulting paper authorship would be mutually beneficial, and provide a comprehensive overview of the task to ensure they make an informed decision to participate. Some of our annotators chose to volunteer as they care deeply about neurodiversity and autism inclusion.
We received IRB approval from our university’s review board and identified volunteer annotators through our association with various academic groups. Given the sensitivity of the content, we provided annotators with appropriate trigger warnings, ensuring they could work at their own pace or withdraw from the study if necessary. Additionally, we connected the members of each annotation team to enable discussions on the content and annotation process as needed.
While our dataset and citations will be made available to the academic community, commercial use of the dataset is not allowed due to the size and nature of the data. As our knowledge of anti-autistic ableism continues to evolve, Autalic’s classification might become outdated. While we will be adding disclaimers if needed to reflect these changes, we still encourage researchers building upon our work to stay updated on the latest semantics by referring to the perspectives of autistic scholars, activists, and organizations. Refer to our Appendix for our Guidelines for Responsible Use.
@@ A.1.7 How to Request Approval
Contact the First Author: All requests and inquiries should be directed to the first author, as listed in the dataset documentation or project website, available here: [REDACTED FOR PRIVACY]
@@ A.6 Self-Agreement Scores
In preliminary studies, we examined different labeling schemes for this task to identify the most efficient and effective annotation strategy. Our experiments revealed high levels of self-disagreement among annotators, as shown in Figure  5. The observed scores (M = -0.06, SD = 0.06) highlight the difficulty of the task and provide a meaningful baseline for comparison. Notably, our own annotation scores for Autalic were higher (M = 0.21, SD = 0.09), suggesting major improvement.
@@ A.7 Annotation Platform
Figure  6 shows an example of an annotation task on our platform with contextual sentences.

### v4

@@ 4.3.3 In-Context Learning
After providing in-context learning examples, Llama3 (+22.96%) and Gemma2 (+12.68%) show the largest relative improvements in F1 scores, indicating that both models benefit from the examples. In particular, our ICL examples also led to substantially more consistent scores across different prompt formulations for each LLM. For instance, supplying Llama3 with examples reduced its relative change in F1 score from 67.49% to 17.40% when switching from PFL to conceptual (AA) prompts, suggesting a better consideration of the connection between anti-autistic ableism and ableism toward autistic people.
@@
Figure 4: Mean Cohen’s Kappa scores for each LLM, comparing agreement with human annotators and other LLMs.
4.3.4 Identifying Ableist and Anti-Autistic Speech
@@
Our results with prompt engineering reveal that anti-autistic ableism is too abstract a concept for LLMs to reliably recognize, providing empirical evidence that they do not adequately reflect the perspectives of autistic people. LLMs struggle to identify anti-autistic speech regardless of the terminology used, further indicating that they are unreliable agents for this data annotation task.
@@
Table 3 presents the results of prompt engineering using either person-first language (PFL), identity-first language (IFL), or the term “anti-autistic” (AA) to frame this form of ableism in more conceptual terms. Notably, switching from PFL or IFL to AA resulted in the largest relative changes in performance. For instance, prompting Llama with “Is this sentence anti-autistic?” instead of “Is this sentence ableist toward people with autism?” led to a relative increase of 67.49% in F1-score. These results indicate that LLMs struggle to frame anti-autistic ableism and ableism toward autistic people as referring to the same phenomenon. Even after introducing ICL examples, the relative change in F1-scores between different prompt types remained high—for example, 32.66% for Gemma2.
@@
Interestingly, DeepSeek has the best results and highest consistency out of all the others. Although its agreement with human annotators is low ($k:0.11$), it is still double that of all other LLMs, as shown in Figure  4. This highlights the difficulty of this task, as more advanced reasoning is required to understand the nuances of anti-autistic ableism, including the terminology we use to describe such speech.
4.4 Error Analysis: Over-classification and Contextual Challenges for LLMs
We analyze the top 10% of sentences in which our human annotators were in perfect agreement, but the LLMs disagreed with them. In doing so, we find many instances where the LLMs over-classified sentences as ableist. For example, Llama labeled 42 sentences as ableist, compared to 0 by our human annotators. Of these 42 sentences, 74-93% were also labeled as ableist by the other LLMs. Notably, 29 of these sentences involved intra-community discussions, highlighting the extent to which community voices may be censored if LLMs are used for content moderation.
While two of the sentences did contain slurs, and 34 included other negatively connoted words—such as “burden”, “threat”, “harm”, or “stupid”—these terms were not used in an anti-autistic context. Figure 5 presents an example of such a sentence. Despite including negatively connoted terms, the sentence expresses an organization’s viewpoint that the author explicitly disagrees with. Since our annotators were instructed to consider context, they correctly labeled this sentence as not ableist. However, the LLMs misclassified it.
This example demonstrates that even when given contextual information, LLMs tend to misclassify sentences based on the mere presence of negative words or connotations. This tendency contributes to their high rates of community censorship if deployed for moderation.
Figure 5: An example of a sentence that LLMs classified as “anti-autistic”, in direct disagreement with all of our human annotators.
@@ 4.5 Discussion
Classifying anti-autistic ableist speech is challenging even within a Western, English-speaking context, as perceptions of what constitutes anti-autistic content differ significantly—even among autistic individuals Keating et al. (2023). Factors such as personal experience, cultural norms, and the evolving discourse around autism advocacy influence each individual’s perception of, and sensitivity to, toxic speech. This variability makes it difficult to establish a consistent classification scheme Kapp et al. (2013); Ousidhoum et al. (2019); Bottema-Beutel et al. (2021b); Taboas et al. (2023).
While developing Autalic, we standardized our definition of anti-autistic ableist speech by providing annotators with an orientation and a glossary. These resources were created in alignment with the perspectives of autistic individuals Bottema-Beutel et al. (2021a); Taboas et al. (2023); Kapp et al. (2013).
Standard baselines such as logistic regression (LR) and BERT performed poorly, underscoring the need for domain-specific fine-tuning. Further, even after fine-tuning BERT on Autalic, our experiments reveal a high false positive rate, which could result in unfair censorship if used for content moderation. Furthermore, our results show that even state-of-the-art LLMs exhibit low agreement with human annotators, highlighting the difficulty of detecting subtle forms of ableism with general-purpose models. Among the models evaluated, DeepSeek achieved the best performance, further emphasizing the complexity of the task.
These findings demonstrate the importance of Autalic in aligning LLM performance with human expectations in the context of autism inclusion and ableist speech classification. Our experiments provide empirical evidence of the current limitations of using LLMs and traditional classifiers to identify anti-autistic ableism. These limitations include misalignment with human perspectives, a lack of conceptual understanding of anti-autistic ableism, and inconsistency—even when provided with in-context learning examples. Each of these challenges hinders the reliability of LLMs for this task.
@@ 5 Conclusion
In this paper, we introduced Autalic, the first dataset specifically focused on anti-autistic ableist language in context. By collecting and annotating 2,400 Reddit sentences along with their surrounding context, we aim to fill a critical gap in current NLP research and improve its alignment with the principles of neurodiversity.
By addressing bias against autistic individuals, Autalic marks a significant step forward in research on ableism within content moderation systems, hate speech detection models, and broader studies of neurodiversity and discrimination. We hope to support the development of more equitable NLP systems that better serve underrepresented and marginalized groups by sharing this resource with the research community,
@@ 6 Limitations
Despite our efforts to include a variety of ableist language targeting autistic individuals, our dataset may still exhibit some selection bias Ousidhoum et al. (2020), as we relied on keyword searches and specific social media threads for data collection.
We acknowledge that, like many datasets presented at ACL, our work primarily reflects Western perspectives Anderson-Chavarria (2022); Thapa et al. (2022); Aoyama et al. (2023); Kirk et al. (2023); Park and Park (2023); Takeshita et al. (2024). Therefore, we do not claim that Autalic is generalizable across languages and cultures, as anti-autistic and ableist speech may manifest and be perceived differently in different contexts. However, we emphasize that Autalic is the first dataset to center the perspectives of autistic people in ableist speech detection, helping to move away from the tendency of AI systems to misclassify any disability-related text as “toxic” Hutchinson et al. (2020); Narayanan Venkit et al. (2023); Van Dorpe et al. (2023); Heung et al. (2024). As such, it can serve as a valuable reference for future work exploring ableist speech in diverse linguistic and cultural contexts.
Our dataset consists of sentences sourced from Reddit, with many posts coming from autism-related subreddits such as r/Aspergers, r/Autism, and r/AutismInWomen. Our search criteria included terms like “r*tard”, a slur broadly used against people with intellectual disabilities, not exclusively targeting autism. While we made efforts to filter out irrelevant posts, our inclusion of broader search terms like “neurodiver” (e.g., “neurodiversity”, “neurodiverse”, and “neurodivergent”) also returned unrelated content, such as discussions about the NEURODIVER video game. Most of the dataset’s content was published in 2023. Although our dataset is relatively small, the quality of the annotations is high, due to standardized protocols and resources developed through rigorous, iterative testing.
@@ 7 Ethics Statement
The data collected for our small-scale, non-commercial research complies with Reddit’s API limits and policies Reddit (2023a, b). All sentences in our dataset are publicly available, and our data collection process follows the methodologies established by prior work Atuhurra and Kamigaito (2024).
We obtained IRB approval from our university’s ethics review board and recruited volunteer annotators through our connections with various academic groups. Given the sensitivity of the content, we provided annotators with appropriate trigger warnings and allowed them to work at their own pace or withdraw from the study at any time. Additionally, we facilitated communication within each annotation team to support discussion of the content and the annotation process as needed. We specifically recruited annotators for whom this collaboration—and potential paper authorship—would be mutually beneficial, and we provided a comprehensive overview of the task to ensure informed consent. Some annotators chose to volunteer due to their strong commitment to neurodiversity and autism inclusion.
While our dataset and associated citations will be made available to the academic community, commercial use—including use by state actors in high-risk domains (e.g., healthcare)—is strictly prohibited. Furthermore, as our understanding of anti-autistic ableism continues to evolve, the classifications in Autalic may become outdated. We will include disclaimers to reflect such developments as needed, but we strongly encourage researchers building on our work to stay informed by consulting current perspectives from autistic scholars, activists, and organizations. Please refer to the Appendix for our Guidelines for Responsible Use.
@@ A.1.7 How to Request Approval
Contact the First Author: All requests and inquiries should be directed to the first author, as listed in the dataset documentation or project website, available here  [^5] https://nrizvi.github.io/AUTALIC.html.
@@ A.6 Self-Agreement Scores
In preliminary studies, we examined different labeling schemes for this task to identify the most efficient and effective annotation strategy. Our experiments revealed high levels of self-disagreement among annotators, as shown in Figure  6. The observed scores (M = -0.06, SD = 0.06) highlight the difficulty of the task and provide a meaningful baseline for comparison. Notably, our own annotation scores for Autalic were higher (M = 0.21, SD = 0.09), suggesting major improvement.
@@ A.7 Annotation Platform
Figure  7 shows an example of an annotation task on our platform with contextual sentences.